What is labor's share of income estimated to be when the only other input to production is physical capital? (α = 1/3)

a. 2/3

b. 0

c. 1/3

d. 1/2
Answer #1

According to the Cobb-Douglas production function -

Capital share of income = α

Labor share of income = 1 - α

α = 1/3

So,

Labor share of income = 1- α = 1- (1/3) = 2/3

Thus,

The labor share of income is estimated to be 2/3.

Hence, the correct answer is the option (a).